OpenStudy (anonymous):

Which equation represents the graph of the linear function? A. y = 4x + 4 B. y = 4x − 4 C. y = −4x + 4 D. y = −4x − 4

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

First, tell me what the y-intercept is.

OpenStudy (igreen):

The y-intercept is the point where the line crosses the y-axis, and where is that?

OpenStudy (anonymous):

4?

OpenStudy (igreen):

No, check again

OpenStudy (anonymous):

is it negative?

OpenStudy (igreen):

Yep, -4.

OpenStudy (igreen):

Now we have the y-intercept. y = mx + b where 'm' is the slope and 'b' is the y-intercept. So far we have: y = mx - 4 Now we find the slope. Take any two points, let's take (0, -4) and (1, 0) and plug them into the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{0+4}{1-0}\) Simplify that

OpenStudy (anonymous):

4/0?

OpenStudy (anonymous):

idk?

OpenStudy (igreen):

4 is correct, but 1 - 0 isn't 0. Check again

OpenStudy (igreen):

Can you subtract 1 - 0? @becker2000

OpenStudy (anonymous):

i mean 1

OpenStudy (anonymous):

@iGreen, Sorry Please come back

OpenStudy (igreen):

Yep, now divide 4 / 1

OpenStudy (anonymous):

4

OpenStudy (anonymous):

So is it B?

OpenStudy (igreen):

Yup, you got it. y = 4x - 4

OpenStudy (anonymous):

Can you help me more? It helps

OpenStudy (igreen):

Sure, just close this one and open a new one.

OpenStudy (anonymous):

ok